OpenStudy (anonymous):

When the equation 3x - 4y = 24 is written in slope-intercept form, what is the value of m?

OpenStudy (anonymous):

i got 3 but just checking, haven't done this in a while. doing the review for the test

OpenStudy (shaik0124):

3x-24=4y y=3/4x-24/4 y=3/4x-6

OpenStudy (anonymous):

so 4??

OpenStudy (shaik0124):

this is slope intercept form y=mx+c

OpenStudy (anonymous):

oh 3/4

OpenStudy (anonymous):

i was kinda confused by the way you wrote it. Thanks for helping! you'll probably be seeing me a lot today...

OpenStudy (shaik0124):

divide both side by 4 then u will get 3/4

OpenStudy (anonymous):

thanks

OpenStudy (shaik0124):

its my pleasure

OpenStudy (anonymous):

How many elements are in the set A ∩ B that means find the ones that are the same correct?

OpenStudy (shaik0124):

for A^B the answer will the similar elements that r in A and B

OpenStudy (anonymous):

ok so heres an example A= 8,9,10 B= 3,6,9 A^B would be 9 right?

OpenStudy (shaik0124):

ya 9 is answer

OpenStudy (anonymous):

thanks

OpenStudy (shaik0124):

kkk

OpenStudy (shaik0124):

wher r u from

OpenStudy (anonymous):

Florida USA

OpenStudy (shaik0124):

kk r u in fbook

OpenStudy (anonymous):

yes why?

OpenStudy (anonymous):

ok whats your name?

OpenStudy (shaik0124):

shaik0124

OpenStudy (shaik0124):

r u der

OpenStudy (anonymous):

yeah and i mean your name on Facebook